# MISTRAL: a model for radiatively efficient AGN winds in cosmological simulations

### v1

@@ Acknowledgements
This work benefited from the "Learning the Universe" collaboration, supported by the Simons Foundation. The simulations presented in this paper have been partly performed using the facilities of the Scientific IT and Application Support Center (SCITAS) of EPFL, using the cluster Jed. This work was also supported by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID s1231. The analysis of the simulations were performed using the Yggdrasil machine, from the Baobab HPC service of the University of Geneva. We are grateful to Volker Springel and the TNG collaboration team for providing access to the arepo code and to the physical modules used in the TNG simulations. We additionally thank Kung-Yi Su for sharing the initial conditions of the idealised galaxy used in this paper, Tiago Costa for his comments on the original draft, and Yohan Dubois for instructive discussions. MF and MH acknowledge funding from the Swiss National Science Foundation (SNF) via a PRIMA Grant PR00P2 193577 “From cosmic dawn to high noon: the role of black holes for young galaxies”. EC was supported by the National Research Foundation of Korea (NRF-RS-2023-00213322). TN acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311 from the DFG Cluster of Excellence "ORIGINS". RW acknowledges funding of a Leibniz Junior Research Group (project number J131/2022).
@@ Appendix A Testing Mistral at higher resolution
Figure 17: Halo mass, stellar mass and black hole mass at $z=3$ for three galaxy simulations at TNG100 (res×$\times 1$) and TNG50 (res×$\times 2$) resolution. Zoom simulations with the fiducial tng physics, mistral-continuous and mistral-stochastic are shown in black, blue and purple, respectively. In each panel, the dashed line shows the 1:1 relation.
@@ Appendix B Calibration of the wind efficiency parameter for Mistral
Our mistral model contains two free parameters: the average wind velocity $v_{\rm w}$, and the AGN momentum feedback efficiency $\epsilon_{\rm w}$. Throughout this paper, we keep $v_{\rm w}=10^{4}\,\rm km\,s^{-1}$, as done by Choi et al. (2012, 2014, 2015, 2017, 2018). In order to determine which $\epsilon_{\rm w}$ should be used in our cosmological zoom simulations, we explore the impact of varying this parameter on 3 of the 15 galaxies from our sample. More specifically, we focus on galaxies hosted in the halos with group number 631, 123 and 57 at $z=2$ (snapshot 33 from TNG100), which correspond to GrNr = 692, 101 and 64 at $z=0$ (snapshot 99 from TNG100, from which their ICs are extracted). We simulate these three halos with mistral-continuous and mistral-stochastic, down to $z=0$, and with $\epsilon_{\rm w}=5\times 10^{-4}$, $10^{-3}$ and $5\times 10^{-3}$.
@@
Figure 18: Scaling relations at $z=0$ for zoom simulations with mistral-stochastic in the upper row (respectively mistral-continuous in the bottom row), with $\epsilon_{\rm w}=5\times 10^{-4},10^{-3},5\times 10^{-3}$ in pink (light blue), purple (blue) and dark purple (dark blue). For reference, we show the counterpart zoom simulations with the fiducial tng physics in black. From left to right: star formation rate and BH mass versus stellar mass, and stellar mass to halo mass relation (displayed at the same x-coordinate). The arrows in the upper left plot denote a SFR lower than the limit of the plot. The legend with the markers indicates the halo group number at $z=2$, for consistency with the rest of the paper.
Figure 19: Same as Fig. 18 but at $z=2$.

### v2

@@ Acknowledgements
We thank the anonymous referee for providing helpful comments which improved this paper. This work benefited from the "Learning the Universe" collaboration, supported by the Simons Foundation. The simulations presented in this paper have been partly performed using the facilities of the Scientific IT and Application Support Center (SCITAS) of EPFL, using the cluster Jed. This work was also supported by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID s1231. The analysis of the simulations were performed using the Yggdrasil machine, from the Baobab HPC service of the University of Geneva. We are grateful to Volker Springel and the TNG collaboration team for providing access to the arepo code and to the physical modules used in the TNG simulations. We additionally thank Kung-Yi Su for sharing the initial conditions of the idealised galaxy used in this paper, Tiago Costa for his comments on the original draft, and Yohan Dubois and Jenny Greene for instructive discussions. MF and MH acknowledge funding from the Swiss National Science Foundation (SNF) via a PRIMA Grant PR00P2 193577 “From cosmic dawn to high noon: the role of black holes for young galaxies”. EC was supported by the National Research Foundation of Korea (NRF-RS-2023-00213322). TN acknowledges support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC-2094 - 390783311 from the DFG Cluster of Excellence "ORIGINS". RW acknowledges funding of a Leibniz Junior Research Group (project number J131/2022).
@@ Appendix A Testing Mistral at higher resolution
[…]
@@ Appendix B Calibration of the wind efficiency parameter for Mistral
Our mistral model contains two free parameters: the average wind velocity $v_{\rm w}$, and the AGN feedback efficiency $\epsilon_{\rm w}$. Throughout this paper, we keep $v_{\rm w}=10^{4}\,\rm km\,s^{-1}$, as done by Choi et al. (2012, 2014, 2015, 2017, 2018). In order to determine which $\epsilon_{\rm w}$ should be used in our cosmological zoom simulations, we explore the impact of varying this parameter on 3 of the 15 galaxies from our sample. More specifically, we focus on galaxies hosted in the halos with group number 631, 123 and 57 at $z=2$ (snapshot 33 from TNG100), which correspond to GrNr = 692, 101 and 64 at $z=0$ (snapshot 99 from TNG100, from which their ICs are extracted). We simulate these three halos with mistral-continuous and mistral-stochastic, down to $z=0$, and with $\epsilon_{\rm w}=5\times 10^{-4}$, $10^{-3}$ and $5\times 10^{-3}$.
@@
Figure 18: Scaling relations at $z=0$ for zoom simulations with mistral-stochastic in the upper row (respectively mistral-continuous in the bottom row), with $\epsilon_{\rm w}=5\times 10^{-4},10^{-3},5\times 10^{-3}$ in pink (light blue), purple (blue) and dark purple (dark blue). For reference, we show the counterpart zoom simulations with the fiducial tng physics in black. From left to right: star formation rate and BH mass versus stellar mass (with observational relations from Kormendy & Ho, 2013; Reines & Volonteri, 2015; Greene et al., 2020), and stellar mass to halo mass relation (displayed at the same x-coordinate, compared with the empirical relation from Behroozi et al., 2019). The arrows in the upper left plot denote a SFR lower than the limit of the plot. The legend with the markers indicates the halo group number at $z=2$, for consistency with the rest of the paper.
Figure 19: Same as Fig. 18 but at $z=2$, with observational estimates from Suh et al. (2020); Abuter et al. (2024); Newman et al. (2025) in the middle panels and from Shuntov et al. (2022) in the right ones.